At a certain temperature, the equilibrium constant, Ke, for this reaction is 53.3. H,(g) +12(R) 근 2H1(R) K, 53.3 At this temperature, 0.700 mol of H2 and 0.700 mol of I2 were placed in a 1.00-L container to react. What concentration of Hl is present at equilibrium? Number
Answer #1

since volume is 1 L , initial concentration will be same as initial number of moles

Let's prepare the ICE table

[H2] [I2] [HI]

initial 0.7 0.7 0

change -1x -1x +2x

equilibrium 0.7-1x 0.7-1x +2x

Equilibrium constant expression is

Kc = [HI]^2/[H2]*[I2]

53.3 = (2*x)^2/(0.7-1*x)^2

sqrt(53.3) = (2*x)/(0.7-1*x)

7.3007 = (2*x)/(0.7-1*x)

5.11048-7.30068*x = 2*x

5.11048-9.30068*x = 0

x = 0.549

At equilibrium:

[HI] = +2x = +2* 0.549 = 1.10 M

Answer: 1.10 M